OpenStudy (anonymous):

write an equation in slope intercept form for the line that passes through the given point and is perpendicular to the given line. (5,-1) y=

OpenStudy (anonymous):

y=4x+7 added to it

OpenStudy (anonymous):

Perpendicular to y=4x+7 means the line you are looking for is at a right angle to this line and therefore has a slope that is the negative reciprocal...that is the slope of the line you are looking for is - 1/4. Now that you have the slope and a point (5,-1) , put it in point slope form first....y-(-1)=-1/4(x-5) and solve for y to get it in slope intercept. your answer should be y=-1/4x+1/4

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

Always remember that a line parallel has the same slope and a line perpendicular has a slope that is the negative reciprocal......:)